OpenStudy (anonymous):

Solve for x. 7x - = 6x -

OpenStudy (anonymous):

thats not the question ops

OpenStudy (anonymous):

\[7x-\frac{ 3 }{ 8 }= 6x - \frac{ 5 }{ 8 }\]

OpenStudy (astrophysics):

Don't let the fractions scare you, do the same thing as before

OpenStudy (astrophysics):

Collect like terms, everything with x on one side and the ones without on the other as follow \[7x-6x = -\frac{ 5 }{ 8 } + \frac{ 3 }{ 8 }\]

OpenStudy (anonymous):

is it subtraction or are they negative fractions that always confuses me @Astrophysics

OpenStudy (astrophysics):

Does what I said above make sense?

OpenStudy (anonymous):

yes, so would it be x=1?

OpenStudy (anonymous):

or 1/4?

OpenStudy (astrophysics):

It is -1/4

OpenStudy (anonymous):

negatives confuse me

OpenStudy (astrophysics):